# Function-Correcting Partition codes

### v1

@@ VI Function Class Privacy in FCPCs
In this section, we discuss a conceptual advantage of FCPCs over FCCs, namely a form of function class privacy with respect to the function being computed. Here, a function class refers to the collection of all functions on the domain $\mathbb{F}_{q}^{k}$ that induce the same domain partition. This notion of privacy arises naturally from the fact that an FCPC depends only on the domain partition induced by the function, and not on the function itself.
@@ VI-A Partial privacy via partition-based encoding
VI-B Function class privacy and special function families
For certain families of functions, such as locally binary functions, additional structure can be exploited to obtain particularly simple or efficient code constructions. In such cases, the domain partition may possess properties that enable very small $\mathcal{D}$-codes. As a construction for $2t$-locally binary functions is given in [1] achieving the lower bound $2t$, which only require two vectors in the corresponding $\mathcal{D}$-code.
When only the partition is revealed, the transmitter may not be able to determine whether the partition arises from a specific special function family, especially when the domain $\mathbb{F}_{q}^{k}$ is large. However, this does not affect the existence or correctness of an optimal FCPC for the given partition. If desired, the receiver may reveal limited additional structural information, such as membership of the function in a particular special function family, along with the partition. This information does not disclose the exact function, but enables the transmitter to employ constructions tailored to that family.
It is possible for two distinct functions to induce the same domain partition and to belong to the same special function family, while differing in their precise mappings. From the transmitter’s perspective, such functions are indistinguishable when only the partition and possibly special function family membership are revealed. The same FCPC applies to both functions, and decoding remains correct for the receiver. Here is one such example.
@@ VI-B Function class privacy and special function families
Consider a setting where the receiver reveals to the transmitter a partition $\mathcal{P}$ of $\mathbb{F}_{q}^{15}$ together with the information that the desired function belongs to the family of $2$-locally binary functions. Equivalently, the receiver specifies that the partition $\mathcal{P}$ is a locally $(2t,2)$-bounded partition for $t=1$. The partition is given by
@@
This partition arises as the domain partition of the Hamming weight distribution function $\Delta_{5}$ on $\mathbb{F}_{q}^{15}$. At the same time, the same partition $\mathcal{P}$ is induced by other functions belonging to the same special function family. For instance, consider the function $f:\mathbb{F}_{q}^{15}\to\mathbb{F}_{2}^{3}$ defined by
@@
where the indicator function $\mathbf{1}\{\mathcal{E}\}$ equals $1$ if $\mathcal{E}$ is true and 0 otherwise. The function $f$ is also a $2$-locally binary function and induces the same domain partition $\mathcal{P}$.
@@
From the transmitter’s perspective, the functions $\Delta_{5}$ and $f$ are indistinguishable when only the partition $\mathcal{P}$ and the information that it is locally $(2t,2)$-bounded are revealed. Consequently, the encoding given by Construction 1 applies to both functions, and decoding remains correct for the receiver.
@@ VII Conclusion
In this work, we introduced function-correcting partition codes as a natural generalization of function-correcting codes by shifting the focus from the function values to the domain partition induced by the function. By focusing on the domain partition rather than the function values, we showed that every FCC is a special case of an FCPC and that a single $(\mathcal{P},t)$-encoding can protect several functions simultaneously through the join of their domain partitions. This naturally led to the notion of partition gains, quantifying the savings in terms of redundancy and rate, compared to straightforward solutions. We established general redundancy bounds for such codes, examined the special case of linear functions through intersections of kernels, and introduced the partition graph to characterize when optimal redundancy can be achieved. Overall, the partition-based viewpoint offers a flexible and efficient approach to multi-function error protection.

### v2

@@ VI Function Class Privacy in FCPCs
In this section, we discuss a conceptual advantage of FCPCs over FCCs, namely a form of function class privacy with respect to the function being computed. Here, a function class refers to the collection of all functions on the domain $\mathbb{F}_{q}^{k}$ that induce the same domain partition. This notion of function class privacy arises naturally from the fact that an FCPC depends only on the domain partition induced by the function, and not on the function itself.
@@ VI-A Partial privacy via partition-based encoding
VI-B Function class privacy and special functions
For certain functions, such as locally binary functions, additional structure can be exploited to obtain particularly simple or efficient code constructions. In such cases, the domain partition may possess properties that enable very small $\mathcal{D}$-codes. As a construction for $2t$-locally binary functions is given in [1] achieving the lower bound $2t$, which only require two vectors in the corresponding $\mathcal{D}$-code.
When only the partition is revealed, the transmitter may not be able to determine whether the partition arises from a special function, especially when the domain $\mathbb{F}_{q}^{k}$ is large. However, this does not affect the existence or correctness of an optimal FCPC for a given partition. If desired, the receiver may reveal limited additional structural information about the function along with the partition. This information does not disclose the exact function, but enables the transmitter to employ constructions tailored to that special function.
It is possible for two distinct functions to induce the same domain partition and to exhibit the same structural characteristics, while differing in their precise mappings. From the transmitter’s perspective, such functions are indistinguishable when only the partition and possibly limited structural information are revealed. The same FCPC applies to both functions, and decoding remains correct for the receiver. Here is one such example.
@@ VI-B Function class privacy and special functions
Consider a setting in which the receiver reveals to the transmitter a partition $\mathcal{P}$ of $\mathbb{F}_{q}^{15}$ together with the information that the desired function is a locally $(2t,2)$-bounded function. Equivalently, the receiver specifies that the partition $\mathcal{P}$ is a locally $(2t,2)$-bounded partition for $t=1$. The partition is given by
@@
This partition arises as the domain partition of the Hamming weight distribution function $\Delta_{5}$ on $\mathbb{F}_{q}^{15}$. At the same time, the same partition $\mathcal{P}$ is induced by other locally $(2t,2)$-bounded functions. For instance, consider the function $f:\mathbb{F}_{q}^{15}\to\mathbb{F}_{2}^{3}$ defined by
@@
where the indicator function $\mathbf{1}\{\mathcal{E}\}$ equals $1$ if $\mathcal{E}$ is true and 0 otherwise. The function $f$ is also a locally $(2t,2)$-bounded function and induces the same domain partition $\mathcal{P}$.
@@
From the transmitter’s perspective, the functions $\Delta_{5}$ and $f$ are indistinguishable when only the partition $\mathcal{P}$ and the information that it is locally $(2t,2)$-bounded are revealed. Consequently, the encoding given by Construction 1 applies to both functions, and the decoding remains correct for the receiver.
@@ VII Conclusion
In this work, we introduced function-correcting partition codes as a natural generalization of function-correcting codes by shifting the focus from the function values to the domain partition induced by the function. By focusing on the domain partition rather than the function values, we showed that every FCC is a special case of an FCPC and that a single $(\mathcal{P},t)$-encoding can protect several functions simultaneously through the join of their domain partitions. This naturally led to the notion of partition gains, quantifying the savings in terms of redundancy and rate, compared to straightforward solutions. We established general redundancy bounds for such codes. Then we focus on how one code can be used to protect multiple linear functions through the intersection of their kernels. We also provide explicit FCPC constructions for locally bounded partitions and grouped weight partitions. Further, we introduce the partition graph to characterize when the problem size to obtain optimal redundancy can be reduced. We also discuss the partial privacy offered by this partition based encoding approach. Overall, the partition-based viewpoint offers a flexible and efficient approach to multi-function error protection.